A car has an initial speed of 50 ft/s and an acceleration of - 4 ft/s2. In the first 10s after the acceleration begins, the car travels
Found 2 solutions by Solver92311, amarjeeth123:
Answer by Solver92311(821)





Plug in your numbers and then evaluate


John

My calculator said it, I believe it, that settles it

From
I > Ø

Answer by amarjeeth123(574)

Initial speed u=50ft/s
Acceleration a=-4 ft/s2
Time t=10s
According to kinematics s=ut+(1/2)at^2 where s is the distance travelled
Plugging in the values we get,
s=50(10)+(1/2)(-4)(100)=500-200=300 ft.
Distance traveled in the first 10 sec=300 ft.